OpenStudy (anonymous):

If two dummies are in a car, and the car is drove into a wall, and dummy A has a wide seat belt on and dummy B has a thin seatbelt on, which dummy has more work on them?

OpenStudy (anonymous):

the dummy with the thinner sidebelt..as the surface area for the thinner sidebelt will be less....

OpenStudy (anonymous):

Assuming the seat belt variation is the only difference between A and B, I believe the the force, F=m*a, would be equal. Because of a smaller area on B, the pressure, P=F/A, would be greater. The work, W=F*d, should be equal. The only argument I could see is if you are considering the properties of the seat belt, then the difference in the acceleration may be significant initially.